An investment club invested part of $10,000 at 4% annual interest and the rest at 8%. If the annual income from these investments was $675, how much was invested at 8%?

Answer by checkley79(3341)

.08X+.04(10,000-X)=675
.08X+400-.04X=675
.04X=675-400
.04X=275
X=275/.04
X=$6,875 AMOUNT INVESTED @ 8%.
10,000-6875=$3,125 AMOUNT INVESTED @ 4%.
PROOF:
.08*6875+.04*3125=675
550+125=675
675=675